if the vectors a, b and c form the sides BC, CA and AB respectively of a triangle ABC, then
axb = bxc = cxa =?

Grade:12

2 Answers

sonika p
6 years ago
if the vectors a, b and c form the sides BC, CA and AB respectively of a triangle ABC, then
axb = bxc = cxa = 2* (area of triangle ABC)
Mohd Mujtaba
6 years ago
Given a ,b and c are vectors along sides of triangle. We know a×b=2area of triangle ABC. Similarly b×c=2area of triangle ABC and c×a=2area of triangle ABC so we get 《a×b=b×c=c×a=2 area of triangleABC》. Thank☺☺☺☺☺☺